An 85 kg box rests on a horizontal surface. The coefficient of static friction between the box and the surface is 0.50. A horizontal force 250 N force is applied to the box. What is the magnitude of the static friction force on the box?

A.42N

B.250N

C.16.5N

D.130N

E.0.00N
Answer #1

Forward force F = 250 N

Now Normal = mg = 85*9.8 = 83.3 N

Friction(max) Fs = mu*N = 0.5*83.3= 416.5  N

since Limiting friction is greater than forward push the friction is static . F< Fs

==> friction = forward push = 250 N

option B is correct.